# Particle Gibbs without the Gibbs bit

### v3

@@ Abstract
Exact parameter and trajectory inference in state-space models is typically achieved by one of two methods: particle marginal Metropolis-Hastings (PMMH) or particle Gibbs (PGibbs). PMMH is a pseudo-marginal algorithm which jointly proposes a new trajectory and parameter, and accepts or rejects both at once. PGibbs instead alternates between sampling from the trajectory, using an algorithm known as conditional sequential Monte Carlo (CSMC) and the parameter in a Hastings-within-Gibbs fashion. While particle independent Metropolis Hastings (PIMH), the parameter-free version of PMMH, is known to be statistically worse than CSMC, PGibbs can induce a slow mixing if the parameter and the state trajectory are very correlated. This has made PMMH the method of choice for many practitioners, despite theory and experiments favouring CSMC over PIMH for the parameter-free problem. In this article, we describe a formulation of PGibbs which bypasses the Gibbs step, essentially marginalizing over the trajectory distribution in a fashion similar to PMMH. This is achieved by considering the implementation of a CSMC algortihm for the state-space model integrated over the joint distribution of the current parameter and the parameter proposal. We illustrate the benefits of method on a simple example known to be challenging for PMMH.
@@ 1 Introduction
This article is concerned with the problem of sampling from the joint distribution of models $\pi_{T}(x_{0:T},\theta)\propto p(\theta)\pi_{T}(x_{0:T}\mid\theta)$, defined recursively as
@@ 3.1 The target distribution
$q(\theta^{\prime}\mid\theta)=\int q(\theta^{\prime}\mid u)q(u\mid\theta)\mathrm{d}u$ (5)
for some auxiliary variable $u$. This decomposition is key in the literature on methods for constructing proposals for ensemble MCMC algorithms (Tjelmeland,, 2004). Examples of these include the case where $q(\theta^{\prime}\mid u)$ and $q(u\mid\theta)$ are two halves of a Gaussian random walk proposal: $q(\theta^{\prime}\mid u)=\mathcal{N}(\theta^{\prime}\mid u,\tfrac{\delta}{2}I)$ and $q(u\mid\theta)=\mathcal{N}(u\mid\theta,\tfrac{\delta}{2}I)$, but also include pre-conditioned versions (see, e.g., Titsias and Papaspiliopoulos,, 2018; Corenflos and Finke,, 2024) or could include more complex proposals combining a gradient oracle $g(\theta)\approx\nabla\log\pi(\theta)$ with a gaussian random walk to form a pseudo-Langevin proposal: $q(\theta^{\prime}\mid u)=\mathcal{N}(\theta^{\prime}\mid u,\tfrac{\delta}{2}I)$ and $q(u\mid\theta)=\mathcal{N}(u\mid\theta+\tfrac{\delta}{2}g(\theta),\tfrac{\delta}{2}I)$.
@@
$\begin{split}\pi(u\mid x^{0:T},\theta^{1:M},l)&\propto q(u\mid\theta^{l})\\ \pi(\theta^{-l}\mid u,\theta^{l},x_{0:T})&=\prod_{j\neq l}q(\theta^{j}\mid u)\end{split}$ (7)
@@ 3.2 Marginal CSMC target
Under these notations, our goal is to form a CSMC algorithm for the marginalized distribution $\pi_{T}(x_{0:T}\mid\theta^{1:M},u)=\sum_{l=1}^{M}\pi_{T}(x_{0:T},l\mid\theta^{1:M},u)$ corresponding to sampling from $x_{0:T}$, marginalizing over the choice of parameter $l$. Once a marginal trajectory $x_{0:T}$ has been obtained, we can then sample the parameter $\theta^{1:M}$ from the distribution $\pi_{T}(\theta^{1:M}\mid x_{0:T})$ in closed form, or using a Hastings acceptance step. To do so, form the sequence
@@
Now, the ratio appearing in the definition of $\gamma_{t}(x_{0:t}\mid\theta^{1:M},u)$ can be decomposed as
@@
$\pi_{t}(l\mid x_{0:t-1},\theta^{1:M},u)=\frac{p_{t}(x_{t}\mid x_{t-1},\theta^{l})g_{t}(x_{t-1:t};\theta)\pi_{t-1}(l\mid x_{0:t-1},\theta^{1:M},u)}{\sum_{m=1}^{M}p_{t}(x_{t}\mid x_{t-1},\theta^{m})g_{t}(x_{t-1:t};\theta)\pi_{t-1}(m\mid x_{0:t-1},\theta^{1:M},u)}$ (12)
@@ 3.3 The algorithm
5:Sample $x^{\prime}_{0:T}$ from Algorithm 1 targeting $\pi_{T}(x_{0:T}\mid u,\theta^{1:M})$.
6:Sample $l^{\prime}\in[N]$ with probability $\pi_{T}(l^{\prime}\mid x_{0:T},u,\theta^{1:M})$.
@@
The algorithm is a valid MCMC kernel targeting the joint distribution $\pi_{T}(x_{0:T},\theta)$ as a collapsed Gibbs sampler (and,, 1994) for a marginalization of the auxiliary index $l$. When $N\to\infty$, it is well known that the CSMC kernel becomes to an exact simulation of the target distribution $\pi_{T}(x_{0:T}\mid u,\theta^{1:M})$ (Andrieu et al.,, 2018, Theorem 1). The sampling of $x^{\prime}_{0:T}$ can therefore be understood as being done marginally with respect to the distribution $\pi_{T}(l\mid u,\theta^{1:M})$, hence the name of the algorithm. Informally, as $N\to\infty$, the algorithm recovers the following steps:
@@
In particular, the last two steps can simply be seen as sampling from $\pi_{T}(l^{\prime}\mid u,\theta^{1:M})$, which, in turn, is a Barker acceptance step (Barker,, 1965) for the parameter $l$ if one were able to compute $\pi_{T}(l^{\prime}\mid u,\theta^{1:M})$ in closed form. This is to be compared with the interpretation of PMMH (Andrieu and Roberts,, 2009) as a noisy version of Metropolis–Hastings acceptance step for the same marginalized distribution. Barker’s acceptance step is known to be at worst half as efficient as Metropolis–Hastings acceptance steps (Łatuszyński and Roberts,, 2013), but also more generally applicable (Vats et al.,, 2021).
@@ 3.4 Choice of the proposal mechanisms
    Closed-form prior mixture: A third choice is to set $p_{t}(x_{t}\mid x_{0:t-1},\theta^{1:M},u)\int p_{t}(x_{t}\mid x_{t-1},\theta^{\prime})q(\theta^{\prime}\mid u)\mathrm{d}\theta^{\prime}$, that is to use the transition density $p_{t}(x_{t}\mid x_{t-1},\theta^{l})$ but marginalize over the proposal $q(\theta\mid u)$. This choice is only valid when the transition density $p_{t}(x_{t}\mid x_{t-1},\theta)$ is conjugate to the proposal $q(\theta\mid u)$, and is not always possible.
@@ 4 Empirical illustration
As explained in Chopin and Papaspiliopoulos, (2020, Section 16.5), for a choice of $q(\theta^{\prime}\mid\theta)=\mathcal{N}(\theta^{\prime}\mid\theta,\tau I)$, with $\tau=0.15^{2}$ (calibrated to obtain a 23.4% acceptance rate of the parameter $\theta$ under the idealized Kalman filter-based MCMC chain), PMMH only gets close to the optimal acceptance rate when $N\gg 1$.
@@
Noting that PMMH is roughly twice as computationally efficient as all our m-PGibbs samplers for the same $N$, we see that m-PGibbs explores the parameter space more efficiently than PMMH for smaller values of $N$. Eventually, for larger values of $N$, the acceptance rate of PMMH converges outperforms m-PGibbs: PMMH converges to the ideal acceptance Metropolis acceptance rate of approximately 27%, while m-PGibbs converges to that of Barker’s acceptance, which in this case is roughly 18%: this is to be expected, as m-PGibbs implicitly uses Barker’s acceptance step (see Section 3.3). Still, m-PGibbs is largely more robust to the choice of $N$ than PMMH, which is a key point in favour of m-PGibbs. Finally, we note that we also compared m-PGibbs with a “classical” PGibbs algorithm for the same proposal, sampling the parameter $\theta$ and the trajectory $x_{0:T}$ in two separate steps. The results are not shown here, because PGibbs failed to mix properly, even for large values of $N$.
@@ 5 Conclusion
This article has demonstrated that the PGibbs algorithm can be marginalized over the parameter space, resulting in a new algorithm called m-PGibbs. We argue that if m-PGibbs can be implemented, it should be preferred to PMMH. While backed up by prior works on CSMC, this is a fairly strong statement, which we have so far only demonstrated up empirically in a simple example. Additionally, a caveat to this is the fact that m-PGibbs uses Barker’s acceptance, which is less efficient than Metropolis–Hastings acceptance used in PMMH: provided one is willing to use a lot of particles, PMMH may eventually outperform m-PGibbs. Augmenting the present method with a Metropolis–Hastings acceptance step is likely possible, via a careful consideration of the augmented distributions and proposals, but we have not explored this in detail. Finally, due to the nature of the algorithm, a theoretical analysis of the algorithm in the style of Andrieu et al., (2018); Lee et al., (2020); Karjalainen et al., (2023) is not straightforward and left for future work. Still, we hope that this article further clarifies the picture of particle MCMC methods that has emerged over the past decade: as a building block, CSMC is a better algorithm than PIMH, and one should keep on building on it rather than defaulting to PIMH.
@@ Acknowledgements
The author thanks jetlag for the idea, and the resulting insomnia for the time to write this article. Special thanks also go to Axel Finke for pointing out a couple of mistakes in the first draft of this article, in particular with respect to the proposal mechanism which originally resulted in a biased Markov chain.

### v4

@@ Abstract
Exact parameter and trajectory inference in state-space models is typically achieved by one of two methods: particle marginal Metropolis-Hastings (PMMH) or particle Gibbs (PGibbs). PMMH is a pseudo-marginal algorithm which jointly proposes a new trajectory and parameter, and accepts or rejects both at once. PGibbs instead alternates between sampling from the trajectory, using an algorithm known as conditional sequential Monte Carlo (CSMC) and the parameter in a Hastings-within-Gibbs fashion. While particle independent Metropolis Hastings (PIMH), the parameter-free version of PMMH, is known to be statistically worse than CSMC, PGibbs can induce a slow mixing if the parameter and the state trajectory are very correlated. This has made PMMH the method of choice for many practitioners, despite theory and experiments favouring CSMC over PIMH for the parameter-free problem. Here, we describe a formulation of PGibbs which bypasses the Gibbs step, essentially marginalizing over the trajectory distribution in a fashion similar to PMMH. This is achieved by considering the implementation of a CSMC algortihm for the state-space model integrated over the joint distribution of the current parameter and the parameter proposal. We illustrate the benefits of method on a simple example known to be challenging for PMMH.
@@ 1 Introduction
This article is concerned with the problem of sampling from the joint distribution of models $\pi_{T}(x_{0:T},\theta)\propto p(\theta)\gamma_{T}(x_{0:T}\mid\theta)$, defined recursively as
@@ 3.1 The target distribution
$q(\theta^{\prime}\mid\theta)=\int q_{\theta}(\theta^{\prime}\mid u)q_{u}(u\mid\theta)\mathrm{d}u$ (5)
for some auxiliary variable $u$ and a pair of (possibly different) proposal distributions $q_{\theta}(\theta^{\prime}\mid u)$ and $q_{u}(u\mid\theta)$. In the remainder of the article, we drop the subscripts for the sake of notational simplicity, and write $q(\theta^{\prime}\mid u)$ and $q(u\mid\theta)$ instead. This decomposition is key in the literature on methods for constructing proposals for ensemble MCMC algorithms (Tjelmeland,, 2004). Examples of these include the case where $q(\theta^{\prime}\mid u)$ and $q(u\mid\theta)$ are two halves of a Gaussian random walk proposal: $q(\theta^{\prime}\mid u)=\mathcal{N}(\theta^{\prime}\mid u,\tfrac{\delta}{2}I)$ and $q(u\mid\theta)=\mathcal{N}(u\mid\theta,\tfrac{\delta}{2}I)$, but also include pre-conditioned versions (see, e.g., Titsias and Papaspiliopoulos,, 2018; Corenflos and Finke,, 2024) or could include more complex proposals combining a gradient oracle $g(\theta)\approx\nabla\log\pi(\theta)$ with a gaussian random walk to form a pseudo-Langevin proposal: $q(\theta^{\prime}\mid u)=\mathcal{N}(\theta^{\prime}\mid u,\tfrac{\delta}{2}I)$ and $q(u\mid\theta)=\mathcal{N}(u\mid\theta+\tfrac{\delta}{2}g(\theta),\tfrac{\delta}{2}I)$.
@@
$\begin{split}\pi(u\mid x_{0:T},\theta^{l})&\propto q(u\mid\theta^{l})\\ \pi(\theta^{-l}\mid u,\theta^{l},x_{0:T})&=\prod_{j\neq l}q(\theta^{j}\mid u)\end{split}$ (7)
@@ 3.2 Marginal CSMC target
Under these notations, our goal is to form a CSMC algorithm for the marginalized distribution $\pi_{T}(x_{0:T}\mid\theta^{1:M},u)=\sum_{l=1}^{M}\pi_{T}(x_{0:T},l\mid\theta^{1:M},u)$ corresponding to sampling from $x_{0:T}$, marginalizing over the choice of parameter $l$. Once a marginal trajectory $x_{0:T}$ has been obtained, we can then sample the parameter $\theta^{1:M}$ from the distribution $\pi_{T}(\theta^{1:M}\mid x_{0:T})$ in closed form, or using a Hastings acceptance step. To obtain the trajectory $x_{0:T}$ in the first place, we define the sequence of marginalized distributions
@@
This recovers the marginal distribution $\sum_{l=1}^{M}\pi_{T}(x_{0:T},l\mid\theta^{1:M},u)$ at the terminal time $T$. Now, the ratio appearing in the definition of $\gamma_{t}(x_{0:t}\mid\theta^{1:M},u)$ can be decomposed as
@@
$\pi_{t}(l\mid x_{0:t-1},\theta^{1:M},u)=\frac{p_{t}(x_{t}\mid x_{t-1},\theta^{l})g_{t}(x_{t-1:t};\theta^{l})\pi_{t-1}(l\mid x_{0:t-1},\theta^{1:M},u)}{\sum_{m=1}^{M}p_{t}(x_{t}\mid x_{t-1},\theta^{m})g_{t}(x_{t-1:t};\theta^{m})\pi_{t-1}(m\mid x_{0:t-1},\theta^{1:M},u)}$ (12)
@@ 3.3 The algorithm
5:Sample $x^{\prime}_{0:T}$ from Algorithm 1 targeting $\pi_{T}(x_{0:T}\mid u,\theta^{1:M})$ using a proposal $p_{t}(x_{t}\mid x_{0:t-1},\theta^{1:M},u)$ and the potential function $g_{t}(x_{0:t};\theta^{1:M},u,l^{\prime})$ defined in (11).
6:Sample $l^{\prime}\in[N]$ with probability $\pi_{T}(l^{\prime}\mid x_{0:T},u,\theta^{1:M})$ where $\pi_{T}(l\mid x_{0:T},u,\theta^{1:M})$ is defined as the terminal distribution of the sequence (8).
@@
The algorithm is a valid MCMC kernel targeting the joint distribution $\pi_{T}(x_{0:T},\theta)$ as a collapsed Gibbs sampler (Liu,, 1994) for a marginalization of the auxiliary index $l$. When $N\to\infty$, it is well known that the CSMC kernel becomes to an exact simulation of the target distribution $\pi_{T}(x_{0:T}\mid u,\theta^{1:M})$ (Andrieu et al.,, 2018, Theorem 1). The sampling of $x^{\prime}_{0:T}$ can therefore be understood as being done marginally with respect to the distribution $\pi_{T}(l\mid u,\theta^{1:M})$, hence the name of the algorithm. Under regularity conditions (Chopin and Singh,, 2015; Andrieu et al.,, 2018), as $N\to\infty$, the algorithm recovers the following steps:
@@
In particular, the last two steps taken together can be seen as sampling from $\pi_{T}(l^{\prime}\mid u,\theta^{1:M})$, which, in turn, is a Barker acceptance step (Barker,, 1965) for the parameter $l$ if one were able to compute $\pi_{T}(l^{\prime}\mid u,\theta^{1:M})$ in closed form. This is to be compared with the interpretation of PMMH (Andrieu and Roberts,, 2009) as a noisy version of Metropolis–Hastings acceptance step for the same marginalized distribution. Barker’s acceptance step is known to be, at worst, half as efficient as Metropolis–Hastings acceptance steps (Łatuszyński and Roberts,, 2013), but also more generally applicable (Vats et al.,, 2021). Sharper results on the relative inefficiency of Barker’s acceptance step are available in the literature (Agrawal et al.,, 2023), and suggest (in the high-dimensional limit) a more optimistic efficiency ratio of approximately $2/3$.
@@ 3.4 Choice of the proposal mechanisms
    Closed-form prior mixture: A third choice is to set $p_{t}(x_{t}\mid x_{0:t-1},\theta^{1:M},u)=\int p_{t}(x_{t}\mid x_{t-1},\theta^{\prime})q(\theta^{\prime}\mid u)\mathrm{d}\theta^{\prime}$, that is to use the transition density $p_{t}(x_{t}\mid x_{t-1},\theta^{l})$ but marginalize over the proposal $q(\theta\mid u)$. This choice is only valid when the transition density $p_{t}(x_{t}\mid x_{t-1},\theta)$ is conjugate to the proposal $q(\theta\mid u)$, and is not always possible.
@@ 4 Empirical illustration
As explained in Chopin and Papaspiliopoulos, (2020, Section 16.5), for a choice of $q(\theta^{\prime}\mid\theta)=\mathcal{N}(\theta^{\prime}\mid\theta,\tau I)$, with $\tau=0.15^{2}$ (this was calibrated in Chopin and Papaspiliopoulos,, 2020, to obtain approximately a 23.4% acceptance rate of the parameter $\theta$ under the idealized MCMC chain), PMMH only gets close to the optimal acceptance rate when $N\gg 1$.
@@
Noting that PMMH is roughly twice as computationally efficient as all our m-PGibbs samplers for the same $N$, we see that m-PGibbs explores the parameter space more efficiently than PMMH for smaller values of $N$. A point of note is that the acceptance rate of PMMH will always be lower than the ideal acceptance rate of Metropolis–Hastings (Andrieu and Roberts,, 2009) owing to its pseudomarginal nature, while m-PGibbs is not pseudomarginal and may improve on Barker’s acceptance rate, as it does here.
For small values of $N$, PMMH simply fails to explore the parameter space, which is in line with prior analyzes of pseudomarginal methods (Doucet et al.,, 2015; Sherlock et al.,, 2015). Eventually, for larger values of $N$, the acceptance rate of PMMH converges outperforms m-PGibbs: PMMH converges to the ideal acceptance Metropolis acceptance rate of approximately 27%, while m-PGibbs converges to that of Barker’s acceptance, which in this case is roughly 18%: this is to be expected, as m-PGibbs implicitly uses Barker’s acceptance step (see Section 3.3). Still, m-PGibbs is largely more robust to the choice of $N$ than PMMH, which is a key point in favour of m-PGibbs. Finally, we note that we also compared m-PGibbs with a “classical” PGibbs algorithm for the same proposal, sampling the parameter $\theta$ and the trajectory $x_{0:T}$ in two separate steps. The results are not shown here, because PGibbs failed to mix properly, even for large values of $N$.
@@ 5 Conclusion
This article has demonstrated that the PGibbs algorithm can be marginalized over the parameter space, forming a new algorithm called m-PGibbs. Altogether, the proposed algorithm has a complexity cost of $O(TNM)$ in the number of time steps $T$, the number of state particles $N$ and the number of parameter particles $M$. While, for notational simplicity, we presented the algorithm for a general $M$, in practice we recommend taking $M=2$, resulting in a computational complexity of roughly twice that of PMMH. This is due to the disappointing performance of ensemble MCMC samplers in the static regime (Pozza and Zanella,, 2024).
We argue that if m-PGibbs can be implemented, it should usually be preferred to PMMH. A caveat to this is the fact that m-PGibbs uses Barker’s acceptance, which is less efficient than Metropolis–Hastings acceptance used in PMMH: provided one is willing to use a lot of particles, PMMH may eventually outperform m-PGibbs. Augmenting the present method with a Metropolis–Hastings acceptance step (rather than Barker’s) is likely possible via a careful consideration of the augmented distributions and proposals, but we have not explored this in detail. While backed up by prior works on CSMC, this is a fairly strong statement, which we have so far only demonstrated up empirically in a simple example. However, due to the nature of the algorithm, a theoretical analysis of the algorithm in the style of Andrieu et al., (2018); Lee et al., (2020); Karjalainen et al., (2023) is not straightforward and left for future work: for example, while PMMH can be proven to have a lower acceptance rate than its “idealized” version corresponding to $N\to\infty$ (Andrieu and Roberts,, 2009), there is no reason why this would be true for m-PGibbs (and it does not seem to be so empirically). Still, we hope that this article further clarifies the picture of particle MCMC methods that has emerged over the past decade: as a building block, CSMC is a better algorithm than PIMH, and one should keep on building on it rather than defaulting to PIMH.
@@ Acknowledgements
The author thanks jetlag for the idea, and the resulting insomnia for the time to write this article. Special thanks go Nicolas Chopin and Axel Finke for suggesting improvements to the manuscript, in particular to Axel for pointing out a mistake in the proposal mechanism which originally resulted in a biased Markov chain. The work was supported by UKRI grant EP/Y014650/1, as part of the ERC Synergy project OCEAN.